OpenStudy (sparklestaraa):

Which statement is correct about the slope and intercept of the two functions? Their slopes are equal, but y-intercepts are not equal. Their slopes are not equal, but y-intercepts are equal. Both slopes and y-intercepts are equal. Neither slopes nor y-intercepts are equal.

OpenStudy (sparklestaraa):

@acxbox22

OpenStudy (acxbox22):

first find the equation of both equations

OpenStudy (sparklestaraa):

Function B- f(x)=3 Function A- 2+5*x

OpenStudy (sparklestaraa):

@confluxepic @Directrix

Directrix (directrix):

Put the y with the equation for A: y = 5x + 2 The second equation is not correct.

Directrix (directrix):

Pick two points from the chart for Line B. Place them in this online calculator and crank out the equation. Post what you get. http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

Directrix (directrix):

@Sparklestaraa

OpenStudy (sparklestaraa):

3x-y+5=0 y=3x+5 @Directrix

Directrix (directrix):

A: y = 5x + 2 B: y=3x+5 Yes, that is what I got for B. Now look at the two equations. Do they have the SAME slope? @Sparklestaraa

Directrix (directrix):

Recall the general form: y = m*x + b where m is the slope and b is the y-intercept. @Sparklestaraa

OpenStudy (sparklestaraa):

they have the same slope i think am i right? @DanJS

Directrix (directrix):

No. A: y = 5x + 2 B: y=3x+5 Look at these. Line A has slope 5 and Line B has slope 3. They are not the same.

Directrix (directrix):

Now, what about the intercepts?

OpenStudy (sparklestaraa):

Ohhh i thought it was 5/3 and 3/5 xD

Directrix (directrix):

Nope. In y = mx + b form, just look at the coefficient of x.

Directrix (directrix):

Let's finish this and move on: Now, what about the intercepts? @Sparklestaraa

OpenStudy (sparklestaraa):

the intercepts are different? @Directrix

Directrix (directrix):

That is correct. So, what is the answer to the question?

OpenStudy (sparklestaraa):

A- y=7x? B- y=8x?

Directrix (directrix):

Look at the question you posted. There is no option that looks like that.

Directrix (directrix):

Look at these. Which one describes the two lines we have looked at? Their slopes are equal, but y-intercepts are not equal .Their slopes are not equal, but y-intercepts are equal. Both slopes and y-intercepts are equal. Neither slopes nor y-intercepts are equal.

Directrix (directrix):

Look at what you wrote in this thread: the intercepts are different? @Directrix

Directrix (directrix):

Look at what I wrote: . Line A has slope 5 and Line B has slope 3. They are not the same. Put those together and select the answer.

OpenStudy (sparklestaraa):

ohhhh D ?

Directrix (directrix):

At long last, that is correct.